OpenStudy (anonymous):

How do I give the multiplicity for each zero? Here is the Problem 2(x-5)(x+4)^2. Explain how you got the answer so I understand.

OpenStudy (anonymous):

I would be able to do a problem like this f(x)=x(x+2)(x+4)

OpenStudy (anonymous):

Refer to the attached plot.

OpenStudy (anonymous):

I don't understand. What does that mean

OpenStudy (anonymous):

\[2 (x-5) (x+4) (x+4)=2 x^3+6 x^2-48 x-160 \]Two of the roots are the same, -4 . Other than that I can not help you.

OpenStudy (anonymous):

hmmmmm...

OpenStudy (anonymous):

??? I'll bump the question because I don't get it.

OpenStudy (freckles):

do you see how there is an exponent 2 on the factor (x+4)?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

\[(x+4)^2=0 \text{ will give you x=-4 multplicity 2 }\]

OpenStudy (freckles):

\[(x-5)^1 =0 \text{ will give you x=5 multplicity 1 }\]

OpenStudy (anonymous):

What exactly does that mean.

OpenStudy (freckles):

that we have the same zero twice

OpenStudy (freckles):

because (x+4)^2=0 (x+4)(x+4)=0 set both factors equal to zero x+4=0 or x+4=0 x=-4 or x=-4 do you see how we have the same zero twice, that is, we have the x=-4 twice?

OpenStudy (freckles):

we say the x=-4 occurs twice or you can use the term multiplicity 2

OpenStudy (anonymous):

And that two at the front?

OpenStudy (freckles):

2 doesn't equal 0

OpenStudy (freckles):

the only things that can equal zero are (x+4)^2 or (x-5) factor that you have

OpenStudy (anonymous):

What I mean is what I do with the 2 in 2(x-5)(x+4)^2

OpenStudy (freckles):

and what i mean is 2 doesn't equal 0 2=0 gives us no zeros because 2 will never be 0 \[2(x-5)(x+4)^2=0 \\ 2(x-5)(x+4)(x+4)=0 \\2=0\\x-5=0\\x+4=0\\x+4=0\] 2=0 gives us no zeros because 2 never equals 0 x-5=0 when x=5 x+4=0 when x=-4 x+4=0 when x=-4

OpenStudy (anonymous):

So I ignore the 2???

OpenStudy (freckles):

yes because the factor 2 cannot be zero